Suppose that the market demand curve for mineral water is given as ?=100−10? and marginal cost is fixed at $4. Find the equilibrium price and quantity in each type of different market structure. Show your calculation.  

a) Monopoly, Cournot duopoly, and Stackelberg duopoly

b)Bertrand duopoly (MR is fixed at the level of MC).

c) Perfect competitive market (MR is fixed at the level of MC).
Answer #1

100-10 P=10-0.1R Ono TRP MR冫10-o.gl-ot 282 jm. 10-0118,-0.29

о. 6 0, 34- 0 2 -0.3 文の 10-0.1(Ao)-0.1(30) 10-2-2 0.02 2=30-0.56, Reades

ru ut v 兀 =EG да!